# quality of service 5G fifth generation

### v2

@@
Semi-Supervised Learning Approach for Efficient Resource Allocation with Network Slicing in O-RAN
@@
$\dagger$ School of Electrical and Computer Engineering, College of Engineering, University of Tehran, Tehran, Iran
@@ Abstract
This paper introduces an innovative approach to the resource allocation problem, aiming to coordinate multiple independent x-applications (xAPPs) for network slicing and resource allocation in the Open Radio Access Network (O-RAN). Our approach maximizes the weighted throughput among user equipment (UE) and allocates physical resource blocks (PRBs). We prioritize two service types: enhanced Mobile Broadband and Ultra-Reliable Low-Latency Communication. Two xAPPs have been designed to achieve this: a power control xAPP for each UE and a PRB allocation xAPP. The method consists of a two-part training phase. The first part uses supervised learning with a Variational Autoencoder trained to regress the power transmission, UE association, and PRB allocation decisions, and the second part uses unsupervised learning with a contrastive loss approach to improve the generalization and robustness of the model. We evaluate the performance by comparing its results to those obtained from an exhaustive search and deep Q-network algorithms and reporting performance metrics for the regression task. The results demonstrate the superior efficiency of this approach in different scenarios among the service types, reaffirming its status as a more efficient and effective solution for network slicing problems compared to state-of-the-art methods. This innovative approach not only sets our research apart but also paves the way for exciting future advancements in resource allocation in O-RAN.

### v3

# Generative Artificial Intelligence xApp eXtended Application

@@
GAI
Generative Artificial Intelligence
xApp
eXtended Application
@@
C-RAN
Cloud Radio Access Network
vRAN
Virtualized Radio Access Network
PDCP
Packet Data Convergence Protocol
SDAP
Service Data Adaptation Protocol
@@
mMTC
Massive Machine-Type Communications
FL
Federated Learning
TL
Transfer Learning
NLOS
non-line-of-sight
RB
Resource Block
AMF
Access and Mobility Management Function
UPF
User Plane Function
OPEX
Operational Expenditures
LLM
Large Language Model
MARL
Multi-Agent Reinforcement Learning
FDRL
Federated Deep Reinforcement Learning
GAN
Generative Adversarial Network
DM
Diffusion Model
Generative AI for O-RAN Slicing: A Semi-Supervised Approach with VAE and Contrastive Learning
@@
School of Electrical and Computer Engineering, College of Engineering, University of Tehran, Tehran, Iran
@@ Abstract
This paper introduces a novel generative AI (GAI)-driven, unified semi-supervised learning architecture for optimizing resource allocation and network slicing in O-RAN. Termed Generative Semi-Supervised VAE-Contrastive Learning, our approach maximizes the weighted user equipment (UE) throughput and allocates physical resource blocks (PRBs) to enhance the quality of service for eMBB and URLLC services. The GAI framework utilizes a dedicated xApp for intelligent power control and PRB allocation. This integrated GAI model synergistically combines the generative power of a VAE with contrastive learning to achieve robustness in an end-to-end trainable system. It is a semi-supervised training approach that concurrently optimizes supervised regression of resource allocation decisions (i.e., power, UE association, PRB) and unsupervised contrastive objectives. This intrinsic fusion improves the precision of resource management and model generalization in dynamic mobile networks. We evaluated our GAI methodology against exhaustive search and deep Q-Network algorithms using key performance metrics. Results show our integrated GAI approach offers superior efficiency and effectiveness in various scenarios, presenting a compelling GAI-based solution for critical network slicing and resource management challenges in next-generation O-RAN systems.